OpenStudy (melissa_something):

Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) 25?

OpenStudy (melissa_something):

x -3 x -3 -3 x A number line with a closed circle on negative 3 and shading to the right.

OpenStudy (anonymous):

7 - 9x - x - 12 = 25 -10x - 5 = 25 10x = - 30 x = -3

OpenStudy (anonymous):

\[7 - 9x - (x + 12) 25 \le 0\] or \[7 - 9x - (x + 12) 25 <0\] ?